# A Survey on Self-play Methods in Reinforcement Learning

### v1

@@ III-A Framework Definition
    Some details of Algo. 23 and 4 need to be mentioned. $\bm{\pi_{opp}}$ denotes opponents’ policies. Moreover, in Algo. 2, off-policy RL algorithms [2, 51, 52, 53], typically require a replay buffer to gather samples; however, for simplicity, they are not explicitly included. Furthermore, Line 1 in Algo. 2 suggests that RL may not always yield an ABR that meets specific criteria. For instance, AlphaGo Zero [9] stipulates that the new policy must achieve a win rate exceeding 55 percent against its predecessor. Therefore, validation of the new policy is sometimes required. In addition, it is important to note that Algo. 3 is specifically devised for R-NaD [54], while Algo. 4 is expressly designed for a series based on regret minimization. They will be introduced in detail in Sec. III-C5 and Sec. III-E respectively.
  • $\text{EVAL}(\Pi)$ (Line 5 in Algo. 1): Evaluating the policy population $\Pi$. There are multiple evaluation metrics available to gauge the performance of each policy. The performance matrix is represented as $\mathcal{P}:=\{p_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{2}}$. Moreover, only effective policies will be evaluated to save on computation. Notably, if the MSS produces a constant matrix irrespective of the input, the evaluation step can be skipped to reduce computations.
Algorithm 2 Compute the Oracle in RL
@@
Algorithm 3 Compute the Oracle in Evolution Theory
@@
Algorithm 4 Compute the Oracle in Regret Matching
@@
Moving forward, we categorize self-play algorithms into four primary groups. We provide a detailed analysis of how each algorithm in these categories integrates and aligns with our proposed framework.
III-B Traditional Self-play Algorithms
Traditional self-play algorithms involve agents improving their strategies by repeatedly playing against themselves, allowing them to explore various strategies and enhance their decision-making abilities without external input. These algorithms can start with agents training against their most recent version, helping to identify and exploit weaknesses. Additionally, other approaches involve training against a set of strategies from different iterations, enabling agents to develop robust and adaptive strategies. In this section, we will explain how traditional self-play algorithms fit into our framework and introduce representative traditional self-play methods, ranging from simpler forms to more complex ones.
III-B1 Integration into Our Framework
We can incorporate the traditional self-play algorithms into our proposed framework (Algo. 1), using the following settings. First, the policy population $\Pi$ utilizes placeholder initialization, meaning that initially, the policies are placeholders rather than actually initialized strategies. This initialization approach is used because the policy population in traditional self-play algorithms is intended to grow with each iteration. Second, we set $E=1$ because in traditional self-play algorithms, only an ineffective policy is likely to be chosen for training in each iteration, thereby turning it into an effective policy. Here, the policy population size $N$ serves as the upper limit for the number of effective policies in the population. In other words, we use $N$ iterations to optimize the policy. Third, the strategy being trained $\pi_{i}^{h}$ can be initialized in a general manner. For instance, the strategy can be initiated randomly, which signifies starting the learning process from scratch. More often, $\pi_{i}^{h}$ is initialized by $\pi_{i-1}(\cdot|h(i-1))$, allowing incremental learning and adaptation based on the most current trained policy to accelerate convergence. Fourth, since the policies in traditional self-play algorithms are not conditioned, we simply set $h(i)=\emptyset$.
Next, we outline the traditional self-play schemes. For the sake of simplicity, we operate under the following assumption:
@@ III-B1 Integration into Our Framework
The policy population gradually increases over time. Consequently, when policy $i$ is selected for training, only policy $j$ $(j\leq i)$ has already been trained and holds meaningful outcomes. Other policies are ineffective policies. As a result, we exclusively select policy $j$, where $j\leq i$, to serve as the opponent for policy $i$. ∎
III-B2 Vanilla Self-play
In vanilla self-play [6], agents are trained by competing against their latest versions, ensuring consistent and aligned learning. The MSS of vanilla self-play:
$\mathcal{F}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}1,\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ j=i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right..$ (33)
No matter what $\mathcal{P}$ is, the MSS produces the same interaction matrix, similar to the iterative refinement process of the policy. Although this MSS is simple, it is utilized in many further works, so we refer to it as vanilla MSS. Although vanilla self-play is effective in transitive games, it can lead the agent to cyclic learning patterns in non-transitive games like Rock-Paper-Scissors.
III-B3 Fictitious Self-play
Fictitious Play (FP) [55] is a learning algorithm in game theory where each player best responds to the empirical frequency of the strategies used by the opponent. If the opponent’s strategy is static, FP can find the NE. Based on FP intuition, Fictitious Self-play (FSP) [13] is introduced to make agents play against past versions of themselves to learn optimal strategies to improve the robustness of vanilla self-play. Neural Fictitious Self-play (NFSP) [56] is a modern variant that combines FSP with deep learning techniques. It uses neural networks to approximate the BRs. In the original versions of NFSP and FSP, two distinct types of agent memory are used: one to record the agent’s own behavior and the other to capture the opponent’s behavior. However, in more recent approaches [42], random sampling is frequently employed to approximate the opponents’ average strategy, eliminating the need to maintain two separate types of agent memory. Thus, the MSS of FSP:
$\mathcal{F}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{i-1},\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ j\leq i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right..$ (34)
@@ III-B3 Fictitious Self-play
III-B4 $\delta$-uniform Self-play
$\delta$-uniform self-play is introduced by [7]. The hyper-parameter $\delta$ ranges between 00 and $1$ used to select the most recent $\delta$ (percentage) of policies for uniform sampling to generate opponent policies. When policy $i$ is in the training phase, following Corollary 1, only the preceding $i-1$ policies hold significance. As opponents for policy $i$, we select policies from the discrete uniform distribution $[\delta(i-1),i-1]$. When $\delta=0$, the system retains the complete historical memory, whereas $\delta=1$ implies that only the most recent policy is utilized. Thus, we can get the MSS of $\delta$-uniform self-play:
$\mathcal{F}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{f(i)},\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \delta(i-1)\leq j\leq i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right.,$ (35)
$f(i)=\max\{\lceil(1-\delta)(i-1)\rceil,1\},$ (36)
@@ III-B4 $\delta$-uniform Self-play
III-B5 Prioritized Fictitious Self-play
Prioritized Fictitious Self-play (PFSP)  [15] leverages a priority function to allocate a higher probability of selection to agents with higher priorities. Here, $\mathcal{P}$ represents the winning rates, specifically defined as $\mathcal{P}_{ij}=\text{Prob}(\pi_{i}\leavevmode\nobreak\ \text{beats}\leavevmode\nobreak\ \pi_{j})$. The MSS of PFSP is given by Algo. 5.
Algorithm 5 the PFSP Meta-strategy Solver
@@ III-B5 Prioritized Fictitious Self-play
The function $f:[0,1]\rightarrow[0,\infty)$ is a priority function. For example, $f(x)=(1-x)^{p}$ with $p>0$ indicates that stronger policies against the currently being trained policy have a higher chance of being chosen as opponents. Alternatively, $f=x(1-x)$ implies that players of similar levels are more likely to be chosen as opponents. Additionally, in broader terms, $\mathcal{P}$ can be assessed by other metrics as well. A similar MSS in  [14] can be utilized to allocate higher probabilities to strategies that perform better or are harder to defeat.
III-B6 Independent RL
Independent RL is a fully decentralized method in MARL. This simplifies the learning process and is useful in competitive or adversarial settings, while it may lead to suboptimal outcomes in situations requiring cooperation. Independent RL can be seen as a special case of self-play algorithms. In each iteration, the policy to be trained, $\pi_{i}^{h}$, is initialized using the previous policy $\pi_{i-1}(\cdot|h(i-1))$. The MSS of independent RL simplifies to an identity matrix:
$\mathcal{F}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}1,\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ i=j\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right..$ (37)
The difference between independent RL and vanilla self-play lies in how the opponents’ policies are handled during training. In vanilla self-play, the opponents’ policies are often fixed, making the training process stationary. In contrast, in independent RL, the opponents’ policies evolve along with the policy being trained, resulting in a non-stationary training process. Moreover, if off-policy RL methods are used in independent RL, the samples collected can still be useful for training, even if they are generated by a different policy. This allows the agent to leverage past experiences more effectively and learn from a wider range of scenarios.
III-C PSRO Series of Algorithms
Similar to traditional self-play algorithms, the PSRO series of algorithms starts with a single policy and gradually expands the policy space by incorporating new oracles. These oracles are policies that approximate optimal responses to the current meta-strategies of other agents. Additionally, PSRO employs EGTA to update meta-strategy distributions, thereby incorporating a level of exploration in policy selection to mitigate the risk of overfitting.
III-C1 Integration into Our Framework
The PSRO series of algorithms can also be integrated into our proposed framework (Algo. 1). First, similar to traditional self-play algorithms, we also utilize placeholder initialization to initialize $\Pi$. Second, we also set $E=1$ and $N$ can be considered as the upper limit for the policy population size in the original PSRO algorithms. Third, in the context of the PSRO series of algorithms, the strategy of our player $\pi_{i}^{h}$ can also be initialized in a general manner. Fourth, we simply set $h(i)=\emptyset$ since the PSRO series of algorithms do not use any conditioning function for their policies. Fifth, it’s crucial to highlight that our framework diverges from the traditional PSRO model [42] in how $\sigma_{i}$ is defined. In contrast to being the meta-strategy for policy, in our framework, $\sigma_{i}$ is the opponent sampling strategy. It means that $\sigma_{i}$ here represents the opponent’s meta-strategy against policy $i$ for the PSRO series of algorithms. Sixth, compared with traditional self-play methods, the MSSes of the PSRO series are often more complex. For example, some MSSes incorporate concepts from different types of game equilibria [45, 46, 47].
For simplicity, we also follow Assumption 1. Similar to traditional self-play algorithms, we can derive the Corollary 2 using a similar proof as Corollary 1.
@@ III-C1 Integration into Our Framework
III-C2 Double Oracle
Double oracle (DO) [45] is traditionally only applied to two-player normal-form games. In this context, we can utilize the payoff matrix as the evaluation matrix. The interaction matrix can be initialized with all zeros, reflecting the initial absence of interactions between strategies. The MSS of DO can then be outlined as described in Algo. 6. The opponent sampling strategy $\sigma_{i}$ corresponds to the opponent’s NE strategy of the restricted game. Therefore, the oracle in DO is a BR rather than an ABR, computing the BR against the current NE opponent strategy of the restricted game. In the context of two-player normal-form games, DO theoretically can achieve the NE of the full game.
Algorithm 6 the NE-based Meta-strategy Solver
@@ III-C2 Double Oracle
III-C3 PSRO
@@
PSRO [42] extends DO to more complex games beyond just two-player normal-form games. This approach first introduces the concept of the MSS, which is a broader concept than simply computing NE. The MSS framework allows for a more flexible representation of strategic solutions in various game settings. Many variants of PSRO focus on designing new MSSes to better capture different aspects of strategic play in these more complex games. In the original PSRO framework, the oracle is computed using RL techniques, similar to those described in Algorithm 2. This allows the algorithm to effectively handle large and intricate strategy spaces, making it applicable to a wide range of game scenarios.
III-C4 PSRO Variants
@@
The traditional PSRO algorithm has been the subject of numerous extensions in recent research. Some studies focus on making PSRO more computationally tractable in different scenarios or achieving fast convergence. Through the establishment of a hierarchical structure of RL workers, Pipeline PSRO [57] achieves the parallelization of PSRO and simultaneously provides guarantees of convergence. Efficient PSRO [58] introduces the unrestricted-restricted (URR) game to narrow down the selection of opponent policies, thereby obviating the need for meta-game simulations in traditional PSRO. And, similar to Pipeline PSRO, Efficient PSRO is equipped to solve URR in parallel. In addition, unlike traditional PSRO, which confines the integration of population strategies to the game’s initial state, XDO [59] allows this integration across all information states. It ensures a linear convergence to an approximate NE based on the number of information states, thereby enhancing its tractability in extensive-form games. ODO [60] integrates the no-regret analysis of online learning with the Double Oracle approach to improve both the convergence rate to an NE and the average payoff. Anytime PSRO [61] and Self-play PSRO [62] are designed to incorporate less exploitable policies into the policy population, thereby facilitating faster convergence. Moreover, as the number of agents grows, determining BRs becomes exponentially challenging. Mean-field PSRO [63] has been introduced to address this complexity in mean-field games. In addition, due to the computational intractability of solving NE in multi-player general-sum games [64, 65] and the selection problem in NE [66], Müller et al. put forward $\alpha$-PSRO [46]. Instead of NE, $\alpha$-rank [30], which is unique and polynomial-time solvable in multi-player general-sum games, is introduced in the MSS and a preference-based best response (PBR) oracle is incorporated in the approach. Similar to $\alpha$-PSRO, [47] proposes JPSRO to tackle multi-player general-sum games. It utilizes concepts of CE and coarse correlated equilibrium (CCE) as alternatives to NE to make it computationally tractable in multi-player general-sum games. It also puts forward the CE and CCE-based MSS. Expanding on JPSRO, NeuPL-JPSRO [67] takes advantage of transfer learning. In addition to a shared parameter $\theta$, each strategy in NeuPL-JPSRO is characterized by a strategy embedding vector $v_{i}$. This approach avoids the need for training from scratch in each iteration, thereby enhancing efficiency.
@@
Other research focuses on policy diversification, given that deriving a singular policy in highly transitive games often lacks significance. [48] introduces an open-ended framework in two-player zero-sum games. This framework enhances the diversity of the strategy population and introduces a gamescape, which geometrically represents the latent objectives in games for open-ended learning. The study proposed two algorithms: Nash response $\text{PSRO}_{\text{N}}$ and rectified Nash response $\text{PSRO}_{\text{rN}}$. Both algorithms utilize an asymmetric payoff matrix as their performance evaluation metric. Similarly to DO, they employ the Nash-based MSS (Algo. 6). Compared to $\text{PSRO}_{\text{N}}$, $\text{PSRO}_{\text{rN}}$ incorporates an additional step within the ABR oracle to focus on the opponents they beat or tie and ignore the opponents they lose to. [49] uses determinantal point process [68] to assess diversity and introduces diverse PSRO by incorporating a diversity term into the PSRO oracle. This modification can also be easily implemented in FP and $\alpha$-PSRO as well. Similarly,  [50] formulates behavioral diversity and response diversity and incorporates them into the PSRO oracle. Policy Space Diversity PSRO [69] defines a diversity metric named population exploitability that helps to achieve a full-game NE.
III-C5 R-NaD

### v2

@@ 3.1 Framework Definition
    Some details of Algo. 23 and 4 need to be mentioned. Algo. 2 is a RL method that can be widely used across different categories. Moreover, off-policy RL algorithms typically require a replay buffer to gather samples; however, they are not explicitly included for simplicity. Algo. 3 is specifically utilized by Regularized Nash Dynamics (R-NaD) [105] and Algo. 4 is specifically utilized by the regret-minimization-based series. They will be introduced in detail in Sec. 3.3.5 and Sec. 3.5 respectively.
  • $\text{EVAL}(\Pi)$ (Line 6 in Algo. 1): Evaluating the policy population $\Pi$. There are multiple evaluation metrics available to assess the performance of each policy. The performance matrix is represented as $\mathcal{P}:=\{p_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{2}}$. $p_{i}$ is the performance of policy i, and the dimension $C_{2}$ depends on the evaluation metric used. For instance, $p_{i}$ can be depicted as the relative skill like Elo ratings ($C_{2}=1$) or can be depicted as the payoff tensor ($C_{2}=N^{n-1}$, where n is the number of players). Specially, in two-player symmetric zero-sum games, the expected payoffs can serve as the evaluation metric. In such cases, $\mathcal{P}$ is a square matrix ($C_{2}=N$).
  • $\textit{MSS}:\mathbb{R}^{N\times C_{2}}\rightarrow\mathbb{R}^{N\times C_{1}}$ (Line 7 in Algo. 1): A meta-strategy solver (MSS) MSS takes the performance matrix $\mathcal{P}$ as its input and produces a new interaction matrix $\Sigma$ as its output. If the MSS produces a constant matrix irrespective of the input, the evaluation step (Line 6 in Algo. 1) can be skipped to reduce computations.
    Moreover, if $\Pi$ uses immediate initialization, $\Sigma$ will be initialized (Line 1 in Algo. 1) by first evaluating the performance and then applying the MSS.
Algorithm 2 Compute the Oracle in RL.
@@
Algorithm 3 Compute the Oracle in Evolution Theory.
@@
Algorithm 4 Compute the Oracle in Regret Matching.
@@
3.2 Traditional Self-play Algorithms
Traditional self-play algorithms involve agents improving their strategies by repeatedly playing against themselves, allowing them to explore various strategies and enhance their decision-making abilities without external input. The simplest form involves agents training against their most recent version to identify weaknesses. Other approaches involve training against a set of strategies from different iterations, enabling agents to develop robust and adaptive strategies. This section will explain how traditional self-play algorithms fit into our framework and introduce representative traditional self-play methods, ranging from simpler forms to more complex ones.
3.2.1 Integration into Our Framework
Traditional self-play algorithms can be incorporated into our proposed framework (Algo. 1) with the following settings. First, the policy population $\Pi$ utilizes lazy initialization because the policy population in traditional self-play algorithms is intended to grow with each iteration. Second, we set $E=1$ because only the inactive policy can be trained in each iteration, then turning into an active policy. Here, the policy population size $N$ serves as the upper limit for the number of active policies in the population. In other words, we use $N$ iterations to optimize the policy. Third, the strategy being trained $\pi_{i}^{h}$ can be initialized in a general manner. For instance, the strategy can be initialized randomly, learning from scratch. More often, $\pi_{i}^{h}$ is initialized by $\pi_{i-1}(\cdot|h(i-1))$, allowing incremental learning and adaptation based on the most current trained policy to accelerate convergence. Fourth, the policies in traditional self-play algorithms don’t need supplementary information, we set the policy condition function $h(i)=\emptyset$. Fifth, the MSSes of traditional self-play algorithms are straightforward, often yielding a constant interaction matrix $\Sigma$ that eliminates performance evaluation. Only the MSS for prioritized fictitious self-play (described in Sec. 3.2.5) requires the performance matrix; nonetheless, it does not need to solve for complex game outcomes such as NE.
Next, we outline the traditional self-play schemes. For simplicity, we operate under the following assumption:
@@ 3.2.1 Integration into Our Framework
The policy population gradually increases over time. Consequently, when policy $i$ is selected for training, only policy $j$ $(j\leq i)$ has already been trained and holds meaningful outcomes. Other policies are inactive policies. As a result, we exclusively select policy $j$, where $j\leq i$, to serve as the opponent for policy $i$. ∎
3.2.2 Vanilla Self-play
In vanilla self-play [108], agents are trained by competing against their latest versions. Thus, the MSS of vanilla self-play:
$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}1,\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ j=i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right..$ (5)
Regardless of $\mathcal{P}$, the MSS produces the same interaction matrix. Although this MSS is straightforward, it is utilized in many other algorithms, so we refer to it as vanilla MSS. Vanilla self-play is effective in transitive games, but it can lead the agent to cyclic learning patterns in non-transitive games. It may also lead to an unstationary problem, causing the system to get stuck in a local optimum.
A more modern version of vanilla self-play is to combine it with off-policy RL. In each iteration, the policy to be trained, $\pi_{i}^{h}$, is initialized using the previous policy $\pi_{i-1}(\cdot|h(i-1))$. Typically, each player’s policy is fixed, and samples are collected into a buffer by vanilla self-play. These samples remain useful for training, even if they were generated by an earlier policy. This allows the agent to leverage past experiences and maintain more stationary training. Once the policy is updated, it will be added to the policy population.
3.2.3 Fictitious Self-play
Fictitious Play (FP) [11] is a learning algorithm in game theory where each player best responds to the empirical frequency of the strategies used by the opponent. If the opponent’s strategy is static, FP can find the NE. Based on FP intuition, Fictitious Self-Play (FSP) [50] is introduced to make agents play against past versions of themselves to learn optimal strategies to improve the robustness of vanilla self-play. Neural Fictitious Self-Play (NFSP) [51] is a modern variant that combines FSP with deep learning techniques. It uses neural networks to approximate the BRs. In the original versions of NFSP and FSP, two distinct types of agent memory are used: one to record the agent’s own behavior and the other to capture the opponent’s behavior.
In more recent approaches [73], random sampling is frequently employed to approximate the opponents’ average strategy, eliminating the need to maintain two separate types of agent memory. Thus, the MSS of FSP:
$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{i-1},\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ j\leq i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right..$ (6)
@@ 3.2.3 Fictitious Self-play
3.2.4 $\delta$-uniform Self-play
@@
$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{f(i)},\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \lceil\delta(i-1)\rceil\leq j\leq i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right.,$ (7)
$f(i)=i-\lceil\delta(i-1)\rceil.$ (8)
In $\delta$-uniform self-play, The MSS generates a constant interaction matrix. Specially, if $\delta=0$, it corresponds to FSP, and if $\delta=1$, it corresponds to vanilla self-play.
3.2.5 Prioritized Fictitious Self-play
Prioritized Fictitious Self-Play (PFSP)  [129] leverages a priority function to allocate a higher probability of selection to agents with higher priorities. Here, $\mathcal{P}$ represents the winning rates, specifically defined as $\mathcal{P}_{ij}=\text{Prob}(\pi_{i}\leavevmode\nobreak\ \text{beats}\leavevmode\nobreak\ \pi_{j})$. The MSS of PFSP is given by Algo. 5. The function $f:[0,1]\rightarrow[0,\infty)$ is a priority function. For example, $f(x)=(1-x)^{p}$ with $p>0$ indicates that stronger policies against the currently being trained policy have a higher chance of being chosen as opponents. Alternatively, $f=x(1-x)$ implies that players of similar levels are more likely to be chosen as opponents. Additionally, in broader terms, $\mathcal{P}$ can also be assessed by other metrics. A similar MSS in  [7] can be utilized to allocate higher probabilities to strategies that perform better or are more challenging to defeat.
Algorithm 5 the PFSP Meta-strategy Solver.
@@ 3.2.5 Prioritized Fictitious Self-play
3.3 PSRO Series
Similar to traditional self-play algorithms, the PSRO series of algorithms starts with a single policy and gradually expands the policy space by incorporating new oracles. These oracles are policies that approximate optimal responses to the current meta-strategies of other agents.
3.3.1 Integration into Our Framework
The PSRO series of algorithms can also be integrated into our proposed framework (Algo. 1). First, similar to traditional self-play algorithms, we also utilize lazy initialization to initialize $\Pi$. Second, we also set $E=1$ and $N$ can be considered as the upper limit for the policy population size in the original PSRO algorithms. Third, in the context of the PSRO series of algorithms, the strategy of our player $\pi_{i}^{h}$ can also be initialized in a general manner. Fourth, we set $h(i)=\emptyset$ since the PSRO series of algorithms do not use any policy condition function for their policies. Fifth, it’s crucial to highlight that our framework diverges from the traditional PSRO model [73] in how $\sigma_{i}$ is defined. In contrast to being the meta-strategy for policy, in our framework, $\sigma_{i}$ is the opponent sampling strategy. It means that $\sigma_{i}$ here represents the opponent’s meta-strategy against policy $i$ for the PSRO series of algorithms. Sixth, compared with traditional self-play methods, the MSSes of the PSRO series are often more complex. For example, some MSSes incorporate concepts from different types of game equilibria [92, 96, 86]. Lastly, we also simply follow Assumption 1. We can derive the Corollary 2 using a similar proof as Corollary 1.
@@ 3.3.1 Integration into Our Framework
3.3.2 Double Oracle
Double Oracle (DO) [92] can be only applied to two-player normal-form games. In this context, we can utilize the payoff matrix as the evaluation matrix. The interaction matrix can be initialized with all zeros, reflecting the initial absence of interactions between strategies. The MSS of DO can then be outlined as described in Algo. 6. The opponent sampling strategy $\sigma_{i}$ corresponds to the opponent’s NE strategy of the restricted game. Therefore, the oracle in DO is a BR rather than an ABR, computing the BR against the current NE opponent strategy of the restricted game. In the context of two-player normal-form games, DO theoretically can achieve the NE of the full game.
Algorithm 6 the NE-based Meta-strategy Solver.
@@ 3.3.2 Double Oracle
We take rock-paper-scissors as an example, a symmetric two-player game. The policy condition function $h(i)=\emptyset$. The policy population size $N=4$. The policy population $\Pi$ uses lazy initialization and the interaction matrix $\Sigma$ is initialized with all zeros. Assume that in the first iteration, the policy $\pi_{1}$ is set to "rock". Obviously, the NE-based MSS returns $\sigma_{2}=(1,0,0,0)$. In the second iteration, the BR oracle adds "paper" to $\Pi$ as $\pi_{2}$. Consequently, the NE-based MSS yields $\sigma_{3}=(0,1,0,0)$, since "paper" always beats "rock." In the third iteration, noting that $\sigma_{3}$ selects only "paper" and never "rock," the BR oracle adds "scissors" as $\pi_{3}$. The NE-based MSS then solves the game and returns $\sigma_{4}=(1/3,1/3,1/3,0)$. In the final iteration, the BR oracle adds a mixed strategy that assigns a uniform probability over "rock," "paper," and "scissors" as $\pi_{4}$. The final interaction matrix is:
$\Sigma=\begin{bmatrix}0&0&0&0\\ 1&0&0&0\\ 0&1&0&0\\ \frac{1}{3}&\frac{1}{3}&\frac{1}{3}&0\end{bmatrix}$ (9)
This process demonstrates how the DO algorithm iteratively expands the strategy set within our framework. Also, using DO as an example, we show that, compared with traditional self-play algorithms, the PSRO series employs more sophisticated MSSes to better capture the essence of the game, albeit at a higher computational cost.
3.3.3 PSRO
@@
PSRO [73] extends DO to more complex games beyond just two-player normal-form games. This approach first introduces the concept of the MSS, which is a broader concept than simply computing NE. The MSS framework allows for a more flexible representation of strategic solutions in various game settings. Many variants of PSRO focus on designing new MSSes to better capture different aspects of strategic play in these more complex games. In the original PSRO framework, the oracle is computed using RL techniques, similar to those described in Algorithm 2. This allows the algorithm to effectively handle large and intricate strategy spaces, making it applicable to various game scenarios.
3.3.4 PSRO Variants
@@
The traditional PSRO algorithm has been the subject of numerous extensions in recent research. Some studies focus on making PSRO more computationally tractable in different scenarios or achieving fast convergence. By establishing a hierarchical structure of RL workers, Pipeline PSRO [87] achieves the parallelization of PSRO and simultaneously provides guarantees of convergence. Efficient PSRO [151] introduces the unrestricted-restricted (URR) game to narrow the selection of opponent policies, thereby preventing the need for meta-game simulations in traditional PSRO. Similar to Pipeline PSRO, Efficient PSRO is equipped to solve URR in parallel. In addition, unlike traditional PSRO, which confines the integration of population strategies to the game’s initial state, Extensive-Form Double Oracle (XDO) [89] allows this integration across all information states. It ensures a linear convergence to an approximate NE based on the number of information states, enhancing its tractability in extensive-form games. Online Double Oracle (ODO) [35] integrates the no-regret analysis of online learning with DO to improve both the convergence rate to an NE and the average payoff. Anytime PSRO [90] and Self-play PSRO [88] are designed to incorporate less exploitable policies into the policy population, facilitating faster convergence. Moreover, as the number of agents grows, determining BRs becomes exponentially challenging. Mean-field PSRO [97] has been introduced to address this complexity in mean-field games. In addition, due to the computational intractability of solving NE in multi-player general-sum games [32, 31] and the selection problem in NE [44], Müller et al. put forward $\alpha$-PSRO [96]. Instead of NE, $\alpha$-rank [101], which is unique and polynomial-time solvable in multi-player general-sum games, is introduced in the MSS and a preference-based best response (PBR) oracle is incorporated in the approach. Similar to $\alpha$-PSRO, [86] proposes Joint Policy-Space Response Oracle (JPSRO) to tackle multi-player general-sum games. It utilizes CE and coarse correlated equilibrium (CCE) as alternatives to NE to make it computationally tractable in multi-player general-sum games and designs the CE and CCE-based MSS.
@@
Other research focuses on policy diversification, as deriving a single policy in transitive games is less meaningful. [4] introduces an open-ended framework in two-player zero-sum games. This framework enhances the diversity of the strategy population and introduces a gamescape, which geometrically represents the latent objectives in games for open-ended learning. The study proposed two algorithms: Nash response $\text{PSRO}_{\text{N}}$ and rectified Nash response $\text{PSRO}_{\text{rN}}$. Both algorithms utilize an asymmetric payoff matrix as their performance evaluation metric. Similarly to DO, they employ the Nash-based MSS (Algo. 6). Compared to $\text{PSRO}_{\text{N}}$, $\text{PSRO}_{\text{rN}}$ incorporates an additional step within the ABR oracle to focus on the opponents they beat or tie and ignore the opponents they lose to. [103] uses determinantal point process [69] to assess diversity and introduces diverse PSRO by incorporating a diversity term into the PSRO oracle. This modification can also be easily implemented in FP and $\alpha$-PSRO. Similarly,  [83] introduces the concepts of behavioral diversity and response diversity and incorporates them into the PSRO oracle. Policy Space Diversity PSRO [142] defines a diversity metric named population exploitability that helps to achieve a full-game NE.
3.3.5 R-NaD